OpenStudy (xenesis98):

GIVES MEDALS!!What is the domain of y = cot x? a. (- ∞ ,∞ ) b.(0,π) c.all x≠ π/2+π d.all x≠ π e.all x≠ π/2

OpenStudy (salazarblack):

The domain of the function y=cot(x)=cos(x)sin(x)y=cot(x)=cos(x)sin(x) is all real numbers except the values where sin(x)sin(x) is equal to 00, that is, the values πnπn for all integers nn. The range of the function is all real numbers.

OpenStudy (xenesis98):

so which would that even be

OpenStudy (xenesis98):

would it be e?

OpenStudy (salazarblack):

it should be a because real numbers are all numbers